### v1

# Transformer-Based Fault-Tolerant Control for Fixed-Wing UAVs Using Knowledge Distillation and In-Context Adaptation

Abstract
This study presents a transformer-based approach for fault-tolerant control in fixed-wing Unmanned Aerial Vehicles (UAVs), designed to adapt in real time to dynamic changes caused by structural damage or actuator failures. Unlike traditional Flight Control Systems (FCSs) that rely on classical control theory and struggle under severe alterations in dynamics, our method directly maps outer-loop reference values—altitude, heading, and airspeed—into control commands using the in-context learning and attention mechanisms of transformers, thus bypassing inner-loop controllers and fault-detection layers. Employing a teacher-student knowledge distillation framework, the proposed approach trains a student agent with partial observations by transferring knowledge from a privileged expert agent with full observability, enabling robust performance across diverse failure scenarios. Experimental results demonstrate that our transformer-based controller outperforms industry-standard FCS and state-of-the-art reinforcement learning (RL) methods, maintaining high tracking accuracy and stability in nominal conditions and extreme failure cases, highlighting its potential for enhancing UAV operational safety and reliability.
I Introduction
In recent years, Unmanned Aerial Vehicles (UAVs) have been widely used to perform various applications in complex and critical scenarios, such as search and rescue or autonomous medical transportation. The operational safety and reliability of these aerial robots have become major concerns due to the potential implications of system failures.
Unlike other robotics fields, such as manipulation and humanoid locomotion, where advanced control methods are essential for managing complex joint movements, UAV Flight Control Systems (FCSs) in industry typically rely on classical control techniques for low-level control layers. While modern approaches, like Model Predictive Control (MPC), offer significant advantages for high-level tasks such as trajectory planning and collision avoidance [1, 2], they require precise system models, extensive uncertainty handling, and high computational resources, which often make them impractical for low-level UAV control. The simplicity, reliability, and efficiency of classical control techniques have established them as the preferred choice for UAVs attitude control.
Figure 1: Trajectory comparison between the proposed transformer-based controller (blue) and an industry-standard FCS (red). Figures (a) and (b) illustrate nominal scenario tracking, while (c) and (d) demonstrate the controllers’ responses to semi-wing damage, with the FCS losing control and the proposed method stabilizing the UAV.
However, complex environments and demanding tasks can cause structural damage to the UAV, altering its aerodynamic characteristics and dynamics. Fixed-wing UAVs, in particular, exhibit highly complex, nonlinear dynamics, which can be significantly disrupted if the structure is compromised. Although current FCSs are robust, they struggle to maintain performance when the vehicle dynamics deviate from the original design specifications, sometimes leading to control divergence and catastrophic failure.
Fault-tolerant flight control has become a focal point for safety-critical UAV operations. Typically, fault-tolerant methods rely on fault detection and diagnosis techniques, which identify faults and then adjust controller parameters to account for the new dynamics [3, 4, 5, 6]. This approach is complex, requiring real-time fault identification and parameter adjustment in a highly nonlinear dynamic environment.
Reinforcement Learning (RL) has introduced alternative solutions to the problem [7, 8, 9, 10]. With its ability to handle high-dimensional, nonlinear dynamics, RL holds promise for system fault management. However, RL algorithms are typically designed for Markov Decision Process (MDP) formulations, while fault-tolerant control—with sudden, unobserved changes in dynamics—must be framed as a Partially Observable Markov Decision Process (POMDP), making RL algorithms learning more difficult and possibly leading them to suboptimal performance [11]. Although recent works have shown possible solutions to this problem [11, 12, 13], these add considerable complexity to the algorithms. Most research focuses on actuator faults in multicopter UAVs [4, 9] or on fixed-wing UAV fault tolerance at the inner attitude control loop level, avoiding altitude, heading, and airspeed tracking [7, 8].
In this work, we propose a novel transformer-based fault-tolerant control method to directly map reference points in the outer loop of attitude UAV control—altitude ($h$), heading ($\Psi$), and airspeed ($V_{T}$)—into control-surface and throttle commands, enabling full UAV control without the need for inner control loops, which complicate the system. Our method employs the attention mechanisms and in-context learning capabilities of transformer models to adapt control actions to dynamic changes during inference, thereby eliminating the need for fault detection or identification and parameter adjustments in the Flight Control System (FCS). Our transformer-based controller uses a context window of past UAV states to autonomously detect and adapt to dynamic changes. Fig. 1 shows a comparison between our proposed transformer-based controller (blue trajectory) and an industry-standard FCS (red trajectory). In Fig. 1(a) and Fig. 1(b), both systems track the commanded references in nominal conditions. Figs. 1(c) and 1(d) illustrate the response when the UAV experiences semi-wing damage, showing the FCS losing control and causing a crash, while our method stabilizes the UAV and follows the reference values.
The contributions of this work are as follows:
  • We present a novel learning framework based on teacher-student knowledge distillation for learning fault-tolerant policies in fixed-wing UAVs. In this framework, the teacher agent is trained using RL on privileged environment information to address partial observability limitations, then a student agent is trained without privileged information using the teacher’s interactions with the environment.
  • We design a transformer-based flight controller that utilizes in-context learning to adapt its behavior in real-time as UAV dynamics change due to failures. This controller employs a context of past states to determine actions, eliminating the need for fault detection methods.
  • We conduct a comparative study against state-of-the-art methods, addressing not only actuator faults but also significant structural damage scenarios.
II Related Works
II-A Learning-Based Fault-Tolerant Control of Aerial Vehicles
Due to the complex dynamics and limited controllability, the problem of fault-tolerant flight control remains challenging. Numerous attempts have been made to address this issue, with data-driven methods gaining popularity in recent years.
Many studies focus on multicopters, often addressing actuator faults [6, 9, 14]. These works demonstrate the solvability of this problem, either through pre-trained fault-sensing networks [6] or RL approaches [14], even when multiple motors fail.
For fixed-wing UAVs, while several studies have applied RL methods to flight-attitude control [15, 16], relatively fewer have focused specifically on fault-tolerant control, which must address a broader range of potential faults beyond actuator failures. An early exploration into using RL for low-level fault-tolerant flight control is presented in Ref. [7], where the authors employ an actor-critic architecture within an inner-loop controller. This approach represents an important step towards RL-based resilience in flight control, with a focus on maintaining stability at lower levels of control. In Ref. [8], RL is combined with Genetic Algorithms (GA) to train a population of agents, offering a novel method that introduces added implementation complexity and highlights the potential of hybrid approaches.
II-B Transformers
Transformers, known for their attention mechanism and advantages in sequence modeling, were initially applied in the field of Natural Language Processing (NLP) [17]. Owing to their high-quality global contextual learning and efficient parallel computation, the transformer architecture has emerged as a powerful alternative to traditional sequence prediction methods like Recurrent Neural Networks (RNNs) and Long Short-Term Memory Networks (LSTMs).
With advances in transformers, their application has extended to RL [18, 19]. Researchers have employed transformer architectures in decision-making agents, proposing the Decision Transformer (DT) model, which frames the sequence modeling problem as an action prediction problem based on historical states and reward-to-go sequences [18]. Trained in a supervised fashion, DT seeks to minimize the error between predicted and ground-truth actions by using sequences of past states and reward-to-go. Recent work has shown that DTs can outperform many state-of-the-art offline RL algorithms [18]. In this work, a DT model is trained to serve as an attitude controller for a UAV in failure scenarios, outperforming other state-of-the-art methods.
II-C Knowledge Distillation
Knowledge distillation is a transfer learning (TL) approach that aims to transfer knowledge from a teacher model to a student model [20]. The original concept of knowledge distillation (KD) involves transferring knowledge from a large, complex model (the teacher) to a smaller, simpler model (the student) [21]. This transfer is typically achieved by minimizing the Kullback-Leibler (KL) divergence between the teacher’s and student’s outputs.
Recent work has combined TL with memory components to address partial observability. In Ref. [22], this combination is used to train a student policy to play soccer based on partial visual observations. In Ref. [23], multiple teachers are distilled into a single multitask student. Authors in Refs. [24] and [25] distill knowledge into a transformer-based student policy to manage partial observability in manipulation tasks and humanoid locomotion, respectively. In contrast, in this work, expert knowledge from the privileged agent is distilled into the DT student in a non-privileged manner, enabling it to adapt its behavior in real time to failures affecting UAV dynamics.
III Problem Statement
In this work, we aim to address the problem of fault-tolerant control in fixed-wing UAVs, where, beyond merely avoiding a crash after a failure, the vehicle is still able to track the given reference values. At a high level of flight-attitude control, we consider the problem of tracking specified set-points in altitude ($h_{\rm ref}$), heading ($\Psi_{\rm ref}$), and airspeed ($V_{T_{\rm ref}}$).
We consider faults that significantly alter the aerodynamic characteristics of the UAV, such as a damaged control surface or a damaged wing. These failures primarily affect the stability and control derivatives, leading to substantial changes in the vehicle’s dynamics. Stability and control derivatives define the forces and moments acting on the UAV, commonly expressed as:
$C_{L}$ $=C_{L_{0}}+C_{L_{\alpha}}\alpha+C_{L_{q}}\frac{qc}{2V}+C_{L_{\delta_{e}}}\delta_{e}$ (1)
$C_{D}$ $=C_{D_{0}}+kC_{L}^{2}+C_{D_{\delta_{e}}}\delta_{e}$ (2)
$C_{Y}$ $=C_{Y_{\beta}}\beta+C_{Y_{p}}\frac{pb}{2V}+C_{Y_{r}}\frac{rb}{2V}+C_{Y_{\delta_{a}}}\delta_{a}+C_{Y_{\delta_{r}}}\delta_{r}$ (3)
$C_{l}$ $=C_{l_{0}}+C_{l_{\beta}}\beta+C_{l_{p}}\frac{pb}{2V}+C_{l_{r}}\frac{rb}{2V}+C_{l_{\delta_{a}}}\delta_{a}+C_{l_{\delta_{r}}}\delta_{r}$ (4)
$C_{m}$ $=C_{m_{0}}+C_{m_{\alpha}}\alpha+C_{m_{q}}\frac{qc}{2V}+C_{m_{\delta_{e}}}\delta_{e}$ (5)
$C_{n}$ $=C_{n_{0}}+C_{n_{\beta}}\beta+C_{n_{p}}\frac{pb}{2V}+C_{n_{r}}\frac{rb}{2V}+C_{n_{\delta_{a}}}\delta_{a}+C_{n_{\delta_{r}}}\delta_{r}$ (6)
These equations define the aerodynamic coefficients $C_{L}$, $C_{D}$, $C_{Y}$, $C_{l}$, $C_{m}$, and $C_{n}$, which represent the lift, drag, side-force, roll moment, pitch moment, and yaw moment coefficients, respectively [26]. Each equation is influenced by the UAV’s dynamic variables (e.g., angle of attack $\alpha$, sideslip angle $\beta$, pitch rate $q$, roll rate $p$, yaw rate $r$) and control surface deflections (e.g., elevator $\delta_{e}$, aileron $\delta_{a}$, rudder $\delta_{r}$). In these expressions, $c$ refers to the mean aerodynamic chord, $b$ is the wingspan of the UAV, $k$ is a factor related to induced drag, and $V$ is the UAV’s airspeed. The coefficients are functions of the stability and control derivatives, which are significantly affected by the UAV’s aerodynamic configuration and can be substantially altered in cases of faults or damage.
The aerodynamic forces and moments derived from the stability and control derivatives dictate the time evolution of the state variables, typically represented as a nonlinear, highly coupled dynamical system, $\dot{x}=f(x,u)$.
Current state-of-the-art control algorithms are unable to adapt to the changes in vehicle dynamics once failure occurs. Therefore, faults and damages in the UAV significantly impact the performance of existing flight control laws, sometimes leading to catastrophic outcomes [27].
To overcome these limitations, we propose a learning-based controller leveraging the power of a transformer model. This new model can use a history of the UAV’s state variables to adapt its behavior in context, without updating its weights.
IV Fault-Tolerant Flight Control Via Sequence Modeling
We formulate the flight-control-under-failure conditions as a partially observable Markov decision process (POMDP). The POMDP is represented by the tuple $\langle A,S,O,P,R\rangle$, consisting of actions $a\in A$, states $s\in S$, observations $o\in O$, transition functions $p\in P$, and rewards $r\in R$. At each timestep $t$, the RL agent receives an observation $o_{t}$ and generates an action $a_{t}$. Then the environment is updated based on the transition function $p(s_{t+1}|s_{t},a_{t})$ and assigns a reward to the agent. The agent’s objective is to maximize cumulative discounted rewards $R_{t}=\sum_{t=k}^{H}\gamma^{t}r_{t}$, where $H$, $k$, and $\gamma$ are the horizon length, current timestep, and discount factor, respectively.
Common POMDP tasks involve challenges where observations only partially reflect the underlying state, where different states may appear identical, or where observations are affected by random noise. Key subareas within POMDP formulations in RL include Meta-RL, Robust RL, and Generalization in RL [11]. In these subareas, agents are designed to handle variations in dynamics and unseen test environments by learning adaptive behaviors. For instance, Meta-RL addresses tasks with episode-specific variations in dynamics, while Robust RL focuses on environments with adversarial perturbations, and Generalization in RL emphasizes resilience to out-of-distribution states [11]. In the context of UAV control under damage and failure, the problem aligns with these subareas, as it involves sudden, substantial shifts in dynamics that are not directly observable, requiring the control policy to adapt without explicit knowledge of the altered dynamics.
IV-A Privileged Agent Training through Online Reinforcement Learning
The primary goal of the RL agent is to minimize the error between the actual values and commanded set-points for the three variables that allow complete control of the UAV in space: altitude ($h$), heading ($\Psi$), and airspeed ($V_{T}$).
The agent takes actions to modify the UAV’s dynamics using control surfaces—aileron ($\delta_{a}$), elevator ($\delta_{e}$), and rudder ($\delta_{r}$)—and throttle ($\delta_{T}$). These actions are based on a set of observations, $o\in O$, consisting of tracking errors ($\epsilon_{h}$, $\epsilon_{\Psi}$, $\epsilon_{V_{T}}$), altitude ($h$), aerodynamic angles ($\alpha$, $\beta$), attitude angles ($\phi$, $\theta$), angular velocity ($p,q,r$), linear velocity ($u,v,w$), and linear accelerations ($n_{x},n_{y},n_{z}$). These observations define the UAV’s attitude.
The reward function is designed to track the reference values while maintaining smooth attitude adjustments and minimizing control effort. The total reward is calculated as a weighted sum of these components:
$\text{r}=\lambda_{1}\cdot r_{\text{h}}+\lambda_{2}\cdot r_{\Psi}+\lambda_{3}\cdot r_{V_{T}}+\lambda_{4}\cdot r_{\text{att}}+\lambda_{5}\cdot r_{\delta},$ (7)
where $r_{\text{h}}$, $r_{\Psi}$, and $r_{V_{T}}$ are tracking rewards, $r_{\text{att}}$ encourages a smooth attitude, and $r_{\delta}$ penalizes large control inputs. The weights $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4},\lambda_{5}$ are set to $0.24,0.2,0.16,0.2,0.2$, respectively.
Tracking reward components take the form:
$r_{x}=-1+\exp\left(-k\cdot\left|\frac{\epsilon_{x}}{\lambda_{\text{s}}}\right|\right),$ (8)
where $x$ represents $h$, $\Psi$, or $V_{T}$, and $k$ and $\lambda_{s}$ are shaping parameters. All reward weights and shaping parameters are initially chosen based on domain knowledge and are further tuned through trial and error to achieve the desired performance.
The attitude reward is calculated using angular rates ($p,q,r$) and roll angle ($\phi$) as:
$r_{\text{att}}=-1+\left(\hat{r_{p}}\cdot\hat{r_{q}}\cdot\hat{r_{r}}\cdot\hat{r_{\phi}}\right)^{1/4},$ (9)
while the control reward uses the deviation of the control commands with respect to the previous time step:
$r_{\delta}=-1+\left(\hat{r_{\Delta\delta_{a}}}\cdot\hat{r_{\Delta\delta_{e}}}\cdot\hat{r_{\Delta\delta_{r}}}\cdot\hat{r_{\Delta\delta_{T}}}\right)^{1/4},$ (10)
For expressions (9)–(10), $\hat{r_{x}}$ takes the form:
$\hat{r_{x}}=\exp\left(-\left(\frac{x}{\lambda_{s}}\right)^{2}\right),$ (11)
where $x\in[p,q,r,\phi,\Delta\delta_{a},\Delta\delta_{e},\Delta\delta_{r},\Delta\delta_{T}]$.
Figure 2: Online training process of the RL algorithm using Domain Randomization and privileged information to enhance adaptability and robustness under varying dynamics. The Domain Randomization (DR) block applies transformations to the physical parameters of the environment, modifying the dynamics. The privileged agent receives as input the concatenation of the current physics parameters and the UAV state obtained from the environment. Based on this information, the agent takes actions over the control surfaces ($\delta_{a},\delta_{e},\delta_{r}$) and throttle ($\delta_{T}$).
To address partial observability, a privileged RL agent is trained. By adding the physical parameters that characterize the environment to the agent’s observations, the POMDP problem is transformed into a classical MDP, resulting in a more efficient and robust agent. As illustrated in Fig. 2, Domain Randomization (DR) is applied to the environment’s physical parameters, altering the UAV’s dynamics at the start and throughout each episode. By combining the agent’s partial observations ($o_{t}$) with the environment’s physical parameters at each timestep, forming the complete state ($s_{t}$), the agent learns to operate across varying UAV configurations and dynamics influenced by aerodynamic, center of gravity, and weight changes across episodes.
The modified physical parameters and their randomization ranges are shown in Table I. Each range is selected based on known physical transformations caused by UAV failure or damage. During training, initial and reference flight conditions ($h$, $V_{T}$) are also randomized within the UAV’s flight envelope.
The privileged agent is trained using the DreamerV3 algorithm [12], selected for its efficiency, robustness, and ability to generalize across multiple environments without extensive hyperparameter tuning. DreamerV3 is an online model-based RL algorithm that learns a world model of the environment from past trajectories, then refines a policy through imagination across the latent space learned in the world model encoder.
TABLE I: Domain Randomization Ranges
Parameter Randomization range
$C_{L_{\alpha}}$, $C_{L_{\dot{q}}}$, $C_{L_{\dot{\alpha}}}$ Scaling $\mathcal{U}(0.6,1.0)$
$C_{D_{0}}$, $C_{D_{i}}$ Scaling $\mathcal{U}(1.0,3.0)$
$C_{D_{mach}}$ Scaling $\mathcal{U}(1.0,2.0)$
$C_{D_{\beta}}$ Scaling $\mathcal{U}(1.0,3.0)$
$C_{Y_{\beta}}$ Scaling $\mathcal{U}(0.5,3.0)$
$C_{Y_{\dot{p}}}$, $C_{l_{\beta}}$, $C_{l_{p}}$, $C_{l_{r}}$, $C_{n_{\alpha}}$, $C_{n_{\beta}}$, $C_{n_{\dot{p}}}$ Scaling $\mathcal{U}(0.5,3.0)$
$C_{Y_{\dot{r}}}$ Scaling $\mathcal{U}(0.5,2.0)$
$C_{m_{0}}$ Additive $\mathcal{U}(-0.02,0.001)$
$C_{m_{\alpha}}$ Scaling $\mathcal{U}(0.5,1.5)$
$C_{m_{q}}$ Scaling $\mathcal{U}(1.0,3.0)$
$C_{m_{\dot{\alpha}}}$ Scaling $\mathcal{U}(1.0,1.5)$
$C_{l_{0}}$, $C_{n_{0}}$ Additive $\mathcal{U}(-0.02,0.009)$
$C_{n_{r}}$ Scaling $\mathcal{U}(0.5,1.0)$
$C_{L_{\delta_{e}}}$, $C_{D_{\delta_{e}}}$, $C_{m_{\delta_{e}}}$, $C_{l_{\delta_{a}}}$, $C_{n_{\delta_{a}}}$ Scaling $\mathcal{U}(0.5,1.0)$
$C_{l_{\delta_{r}}}$, $C_{n_{\delta_{r}}}$, $C_{Y_{\delta_{r}}}$ Scaling $\mathcal{U}(0.0,1.0)$
$x_{CG}$, $y_{CG}$, $z_{CG}$ Scaling $\mathcal{U}(0.70,1.30)$
$W$ Scaling $\mathcal{U}(0.70,1.05)$
  • Modification ranges for the physical parameters—stability and control derivatives, center of gravity position, and weight—in the environment during training. Values are periodically changed throughout episodes by sampling from a uniform distribution, with limits for each parameter chosen based on domain knowledge.
IV-B Knowledge Distillation Via Decision Transformer
Figure 3: Knowledge distillation process through offline reinforcement learning using partial observations of the POMDP, derived from expert trajectories generated by the privileged agent.
Building upon the trained expert privileged policy, we employ offline reinforcement learning to train a student policy through knowledge distillation, using only the partial observations of the environment ($o\in O$).
As shown in Fig. 3, the privileged agent is used to collect a dataset of multiple expert trajectories, encompassing random variations in physical parameters and flight conditions. From these trajectories, we retain only the partial observations ($o_{t}$), taken actions ($a_{t}$), and rewards ($r_{t}$). This dataset, consisting of expert trajectories of the MDP reduced to a POMDP of partial observations, is used to train a Decision Transformer (DT) [18] through offline RL.
Decision Transformers are powerful models capable of leveraging in-context learning to adapt behavior without updating weights. Thus, the DT can learn to take optimal actions across different system dynamics using only a history of partial observations and actions. Ultimately, a DT trained on multiple expert trajectories can generalize to unseen failure scenarios and adapt its behavior to changes in the UAV dynamics using only the history of past observations to identify variations.
The final model is a causal decoder-only transformer of size $\sim$0.8M, comprising 4 heads and 3 layers, with an embedding dimension of 128 and a maximum sequence length of 60 steps (6 s). Training is performed with a batch size of 256, a learning rate of 1e-04, and a weight decay of 1e-04. The model was trained on a dataset of 2 million timesteps, encompassing 200,000 trajectories.
V Experiments
In this section we present several experiments to validate the effectiveness of our proposed method. We first provide a description of the testing cases designed to validate our method in specific failure scenarios rather than randomized variations, as well as the implementation details of the experiments. Secondly, we demonstrate our model’s capability through a comparative study across different cases.
V-A Testing Cases and Implementation Details
We evaluate our method across several failure and damage scenarios for the UAV. Unlike the random and independent parameter modifications used during training, testing cases involve modifications to aerodynamic parameters and control surfaces based on the changes that a specific failure produces in the UAV’s dynamics, creating unseen situations for the agent. The defined scenarios are:
  • Jammed Rudder: Rudder stuck at a deflection of 15°.
  • Broken Aileron: Aileron effectiveness decreased by 50% with an increase in the parasitic roll moment due to asymmetric force.
  • Saturated Elevator: Elevator deflection limited to $\pm$5of the maximum range.
  • Deployed Landing Gear: Sudden deployment of landing gear, resulting in increased drag given by $3C_{D_{\text{gear}}}$, which was unseen during training.
  • Shifted Center of Gravity: UAV center of gravity shifts along the $x$, $y$, and $z$ axes by +20%, +40%, and -30%, respectively.
  • Damaged Horizontal Tail: Partial tail damage causing reduced lift ($0.5C_{L_{\delta_{e}}}$), increased drag ($1.2C_{D_{\delta_{e}}}$, $1.2C_{D_{0}}$), and affecting pitch moment ($0.5C_{m_{\delta_{e}}}$, $C_{m_{0}}$, $0.5C_{m_{q}}$, $0.5C_{m_{\alpha}}$, $0.5C_{m_{\dot{\alpha}}}$).
  • Damaged Semi-Wing: Damaged wing producing a reduction in lift effectiveness ($0.7C_{L_{\alpha}}$, $0.7C_{L_{q}}$, $0.7C_{L_{\dot{\alpha}}}$), increased drag ($1.2C_{D_{0}}$, $1.2C_{D_{i}}$), and affected side force ($1.5C_{Y_{\beta}}$, $1.5C_{Y_{p}}$, $1.5C_{Y_{r}}$). Roll moment altered ($C_{l_{0}}$, $-1.0C_{l_{\beta}}$, $-1.0C_{l_{r}}$, $0.5C_{l_{\delta_{a}}}$), and yaw moment modified ($C_{n_{0}}$, $0.8C_{n_{r}}$, $1.5C_{n_{\beta}}$, $1.5C_{n_{\delta_{a}}}$, $1.2C_{n_{\alpha}}$).
During validation, multiple episodes are run in which the UAV’s dynamics are changed from the nominal condition to the modifications defined for each failure scenario at a specified timestep. Set-point changes for ($h,\Psi,V_{T}$) also occur at specified timesteps.
Experiments and training are conducted using JSBSim, a high-fidelity 6-DoF flight dynamics model. The UAV physics model used operates within flight conditions of $h\in[4000,24000]$ ft and $V_{T}\in[260,360]$ ft/s. Episodes have a maximum length of 1000 steps. Light-intensity wind gusts are enabled during validation, following a Gaussian distribution N(0, 2 ft/s) for the north, east, and down gust components.
V-B Comparison Study
TABLE II: Comparison Study Results
Scenarios DT (Ours) RL+DR RL FCS FCS+RL
Nominal $\mu\pm\sigma$124.61 ± 25.25 -232.59 ± 90.27 -205.53 ± 46.71 -257.30 ± 34.15 -259.80 ± 36.42
Crash % 0.00% 0.00% 0.00% 0.00% 0.00%
Jammed Rudder $\mu\pm\sigma$398.49 ± 25.69 -409.40 ± 55.31 -412.40 ± 79.35 -585.60 ± 393.28 -543.04 ± 390.62
Crash % 0.00% 0.00% 0.00% 20.00% 32.00%
Broken Aileron $\mu\pm\sigma$151.94 ± 26.67 -398.75 ± 70.36 -409.06 ± 24.57 -485.19 ± 211.83 -430.50 ± 257.86
Crash % 0.00% 0.00% 0.00% 4.00% 16.00%
Saturated Elevator $\mu\pm\sigma$220.11 ± 74.98 -301.69 ± 54.27 -301.84 ± 69.89 -545.53 ± 35.78 -443.64 ± 31.37
Crash % 0.00% 0.00% 0.00% 0.00% 0.00%
Deployed Landing Gear $\mu\pm\sigma$216.02 ± 73.54 -307.10 ± 77.71 -270.45 ± 27.50 -552.65 ± 306.27 -382.85 ± 92.94
Crash % 0.00% 0.00% 0.00% 12.00% 8.00%
Shifted CG $\mu\pm\sigma$142.78 ± 19.45 -349.35 ± 94.51 -272.70 ± 51.39 -733.69 ± 420.98 -414.10 ± 103.58
Crash % 0.00% 0.00% 0.00% 28.00% 20.00%
Damaged Tail $\mu\pm\sigma$202.10 ± 40.86 -430.47 ± 46.02 -349.36 ± 28.81 -721.62 ± 436.57 -402.62 ± 232.80
Crash % 0.00% 0.00% 0.00% 28.00% 8.00%
Damaged Wing $\mu\pm\sigma$310.98 ± 23.46 -349.55 ± 71.68 -1394.93 ± 155.40 -1167.78 ± 338.98 -762.48 ± 444.57
Crash % 0.00% 0.00% 100.00% 92.00% 60.00%
  • Comparison of performance results across different algorithms. DT refers to our method. RL+DR denotes the DreamerV3 agent trained with Domain Randomization (DR) without privileged environment information. RL refers to the DreamerV3 agent assuming an MDP formulation. FCS represents the baseline industry-like designed FCS. FCS+RL refers to the FCS with added Gain-Scheduler (GS) trained using RL. Performance is measured using the mean and standard deviation of episode returns and the crash percentage across 100 episodes for each failure scenario.
We evaluate our proposed method against several state-of-the-art flight control systems (FCS) to showcase the advantages of our approach in fault-tolerant flight control.
  • Privileged Policy: The teacher policy trained with privileged information of the environment.
  • RL Policy: RL-trained policy without environmental changes, assuming an MDP. This policy is tested in the POMDP setting and trained using DreamerV3.
  • RL Policy + DR: RL policy trained with environment variations through domain randomization, considering the setup as a POMDP.
  • Industry FCS: A baseline FCS following common industry practices, with separate control for lateral-directional and longitudinal motions using nested PID controllers.
  • Industry FCS + RL + DR: A RL policy trained to select the PID gains of the FCS at each step, following a Gain-Scheduling (GS) approach for fault-tolerant control. The policy is trained in the POMDP setting with domain randomization.
Figure 4: Comparison between the privileged agent teacher and DT student policies across each failure scenario, using mean episode return as the performance metric.
Figure 5: Sample trajectories comparing our DT method against the baseline FCS, RL trained with DR (RL+DR), base RL agent (RL), and trained GS system (FCS+RL) in the nominal and damaged wing scenarios for tracking reference values of altitude ($h$), heading ($\Psi$), and airspeed ($V_{T}$). (a) Nominal scenario. (b) Damaged wing scenario. The dashed black line represents the setpoint for each value, and the vertical dashed red line marks the timestep at which the failure occurs.
To assess the effectiveness of the DT under POMDP conditions, we compare the results of the DT student agent and the privileged teacher agent. Both policies are tested across all scenarios in 100 episodes each, with flight conditions and random seeds standardized for fair comparison. Fig. 4 shows the results, using the mean episode return as the performance metric, where the theoretical maximum is zero. Results demonstrate similar performance between the agents, showing the DT’s capability to learn behaviors from the expert dataset under the POMDP setting and to adapt to dynamic variations in context. In some scenarios, the DT even outperforms the teacher agent, indicating that it learns through offline RL rather than simply mimicking the training data through imitation learning (IL).
A full comparison study, as shown in Table II, assesses the DT’s performance against other algorithms, including state-of-the-art online RL agents and industry-standard FCS commonly used in commercial UAVs. Performance is measured by the mean and standard deviation of episode returns, as well as crash percentage due to total control loss. Fatal failures terminate an episode, occurring under extreme values of angular rates or accelerations, or when the UAV breaches lower altitude limits. As before, 100 episodes are run for each failure scenario. Results in Table II show that the DT policy outperforms others in all failure and nominal scenarios, effectively tracking set-points and maintaining control even under severe damage and extreme conditions.
Promising results are also obtained from the RL agent with domain randomization, showing high tracking performance and achieving zero crashes across failure scenarios. These results highlight the potential of model-based RL algorithms with added memory, such as DreamerV3. However, the DT agent’s superior performance, despite its significantly smaller parameter size ($\sim$0.8M) compared to DreamerV3 ($\sim$26M), underscores the potential of transformer-based models for real-time dynamics adaptation through in-context learning.
Figure 6: Angular rates ($p$, $q$, $r$) following wing damage, comparing DT agent and baseline FCS responses. The dashed red line marks the time step at which failure occurs.
Currently used FCSs, as described, are difficult to replace due to their simplicity, reliability, modularity, and computational efficiency, all while avoiding the need for an accurate UAV model and handling uncertainties. However, they lack adaptability to significant dynamic changes. Adding a Gain Scheduling (GS) algorithm to the FCS could enhance system adaptability under failure conditions by modifying gains in real time; however, GS faces challenges related to computational demands and nonlinear dynamics. Even state-of-the-art RL algorithms trained for fault tolerance struggle to achieve high performance across scenarios, as shown in Table II.
Our method takes advantage of the in-context learning capabilities of transformers to adapt policy behavior during flight without changing weights, rather than modifying controller gains. Fig. 5 presents sample trajectories of our method versus the other algorithms in the comparison study for tracking $h$, $\Psi$, and $V_{T}$. In Fig. 5(a), the nominal UAV dynamics are shown, where all designed controllers track reference values with similarly high performance, with some noise in the measured airspeed due to light wind gusts.
Fig. 5(b) shows the trajectory for the damaged wing scenario, the most challenging due to the significant dynamic changes. In this sample, the failure occurs at timestep 100 (indicated by the vertical dashed red line), and the setpoint change requirement occurs at timestep 200. While the Baseline FCS, the base RL agent, and the FCS with added GS (FCS+RL) diverge, losing control and resulting in a crash and episode termination (triggered by either extreme accelerations or entering a tailspin), the RL agent trained with DR (RL+DR) and our DT method regain control and track setpoints effectively despite substantial damage. Among these, the DT method achieves the highest tracking accuracy for all three reference values showing superior performance.
Fig. 6 illustrates angular rates ($p$, $q$, $r$) for our DT method and the baseline FCS in the damaged wing scenario. The vertical red line marks the onset of failure, highlighting the abrupt disturbance caused in the roll rate ($p$). After this point, the controller must regain control of the UAV to re-stabilize it to the previous flight condition. In this case, our DT agent successfully stabilizes the UAV, converging to the prior steady flight state despite the damaged aerodynamics and control surfaces. In contrast, the baseline industry-designed FCS begins to diverge due to accumulating errors over time, eventually leading to total control loss and stall.
VI Conclusion
In this paper we have introduced a novel transformer-based approach for fault-tolerant control in fixed-wing UAVs, leveraging the transformer’s in-context learning capabilities to adapt in real-time to sudden changes in vehicle dynamics caused by structural and actuator failures. Unlike traditional FCSs, which struggle to maintain performance under substantial dynamic deviations, our method directly maps outer-loop reference values to control commands, bypassing the need for inner-loop controllers and additional fault detection mechanisms.
Through a teacher-student knowledge distillation framework, our approach trains a student policy on partial observations using knowledge transferred from a privileged expert agent. This enables the student to generalize across a range of complex failure scenarios, including actuator faults and structural damage, achieving high performance without requiring privileged environment information.
Experimental results demonstrate that our transformer-based control method not only outperforms industry-standard FCS and state-of-the-art reinforcement learning approaches but also remains resilient across challenging test cases. By effectively utilizing the transformer’s attention mechanism, our approach exhibits robust adaptability, maintaining control even in the face of severe aerodynamic and structural disruptions.
Future work will explore further enhancements to the model’s generalization capabilities by incorporating more complex training environments and testing on a wider array of fault conditions. Additionally, extending this transformer-based approach to other types of autonomous vehicles would help advance fault-tolerant control across various domains.